OpenStudy (anonymous):

what is the point of discontinuity for y=x^2-x-12/-x^2-2x+3

OpenStudy (tkhunny):

x = 0 might be. Of course, if you actually mean y=(x^2-x-12)/(-x^2-2x+3), there will be a different answer. Please remember your Order of Operations. Include necessary parentheses.

OpenStudy (jdoe0001):

\(\large { y=\cfrac{x^2-x-12}{{\color{brown}{ -x^2-2x+3}}}\quad if\quad {\color{brown}{ -x^2-2x+3}}=0\quad \\ \quad \\ then\quad y=\cfrac{x^2-x-12}{{\color{brown}{ 0}}}\leftarrow \begin{cases} undefined\\ discontinous \end{cases} }\) so see what values of "x" make the denominator 0